# The Hierarchical Growth of Bright Central Galaxies and Intracluster Light as Traced by the Magnitude Gap

### v1

@@ 4 The Statistical Modelling of the SMHM Relation
where is the mathematical offset, is the slope of the SMHM relation, is the M14 parameter, and is the intrinsic scatter in stellar mass at fixed . This is a nested model, if =0, Equation 4 reverts to Equation 5. The only differences between the model used in this analysis and those in Golden-Marx & Miller (2019) and Golden-Marx et al. (2023) are that we do not account for redshift evolution and we include , respectively.
A Bayesian formalism works by convolving prior information associated with a selected model with the likelihood of the observations given that model, to yield the posterior distribution for the non-nuisance parameters in our model. To determine the posterior distributions for each SMHM parameter, our MCMC model generates values for the observed aperture stellar masses, richness-estimated halo masses, and M14s at each step in our likelihood analysis, which are directly compared to the observed measurements.
We model the log10 BCG stellar masses (), log10 halo masses (), and M14 values () as normal distributions with mean values (locations) taken from our measurements. The standard deviations associated with each value is taken from measurement uncertainties, described in Section 3 (, , ) as well as a stochastic component from a beta function, (Golden-Marx & Miller, 2018), which allows for additional uncertainty on the observational errors. These errors are treated statistically in the Bayesian model as free nuisance parameters , , and . Additionally, as done in Golden-Marx & Miller (2019), Golden-Marx et al. (2022), and Golden-Marx et al. (2023), we reduce the covariance between and by subtracting off the median values of the upper and lower limits in log/ and log/, 11.3 and 14.55 respectively. For consistency we subtract off the median stellar mass measured within 0-30 kpc in each aperture bin.
@@
The parameters of the SMHM relation are given in Table 1, where each cluster is a component in the summed log likelihood and the terms marked are representative of the observed data.
@@ 5 The Observed SMHM Relation
As discussed in Section 1, many recent analyses have focused on the connection between the ICL and the cluster’s dark matter halo. Recent observations (e.g., Montes & Trujillo, 2019) suggest the light distribution of the ICL may trace the halo’s dark matter distribution. Moreover, prior results have identified statistically significant correlations between the stellar mass contained within the ICL and the cluster’s halo mass (e.g., Sampaio-Santos et al., 2021; Huang et al., 2022; Golden-Marx et al., 2023). As presented in Golden-Marx et al. (2023), accounting for the ICL increases the slope, , of the SMHM relation by 0.1-0.15. Thus, as a result of BCG+ICL hierarchical growth, the larger slope is measured because ICL, due to recent growth, is more correlated with the dark matter halo mass. Here, for the first time, we analyse the correlation between M14 and the ICL and investigate both quantitatively and qualitatively the impact of incorporating the ICL and M14 into the SMHM relation.
@@ 5.1 The BCG+ICL+M14 SMHM relation
Figure 1 visualizes the stellar mass within the selected aperture plotted against the halo mass for each annulus, colour coded by M14. In each subplot, we overplot three lines, which are representative of the median of one-third of the data sorted by M14. The median M14 values are 0.90 (dark blue), 1.50 (green) and 2.10 (yellow). In the final subplot, we overlay the posterior distribution from our MCMC analysis for samples representative of the upper and lower M14 values to illustrate our fitting.
@@
To quantify the impact of including the ICL and M14 on the parameters of the SMHM relation, we present the values of the posterior distribution (both for when M14 is and is not incorporated) in Table 3. For consistency, we see that the results in the radial range of kpc are in excellent agreement with the results presented in Golden-Marx & Miller (2019) and Golden-Marx et al. (2022). Moreover, these SMHM relation parameters are within 1 of the values of and measured in Golden-Marx et al. (2023) (due to the different median subtraction values, we can’t compare ), which shows our results are not biased by the use of richness-inferred halo masses as opposed to SZ-estimated halo masses.
@@
Across each aperture, we see clear trends such that the incorporation of M14 increases by 0.10-0.15 matching prior results (Golden-Marx & Miller, 2018, 2019; Golden-Marx et al., 2022). Moreover, reaches an asymptote at a radii of 150 kpc, in strong agreement with the trends found in Golden-Marx & Miller (2019) and Golden-Marx et al. (2023). Additionally, the inclusion of the ICL increases by 0.1-0.15 as found in Golden-Marx et al. (2023) (a similar increase to what is found by including M14). As in Golden-Marx & Miller (2019), the value asymptotes at is significantly higher when M14 is accounted for, by approximately 0.1.
@@
Additionally, although Figure 2 does not present the measurements, the values given in Table 3 are all greater than 13 from 0. Thus, the M14-stellar mass correlation persists out to 300 kpc. However, decreases as we progress to larger radii, which can be seen by the separation between the solid lines shown in Figure 1.
@@
The presence of a non-zero along with the reduction in and increase in allow us to conclude that M14 is a latent parameter when the SMHM relation is measured out to the ICL. For further verification, we randomly reassign the M14 values. If a non-zero value is measured for , this would suggest an intrinsic, non-physical, correlation between these parameters. However, as shown in Table 6, for each bin, the shuffled value is equivalent to 0. Moreover, the values of , , and are in excellent agreement between the versions with the randomly assigned M14 (Table 6) and when M14 is not incorporated (Table 3). Thus, any change in parameter measurements results solely from a correlation between the BCG+ICL stellar mass and M14. Going forward, we include M14 in all measurements discussed in this analysis (though we provide the posteriors when M14 is not accounted for in Tables 3, 4, and 5).
5.2 Comparing the SMHM relation of the BCG to the SMHM relation of outskirts
@@
Although the correlation between the stellar mass and the halo mass increases as the BCG+ICL light profile is extended outward, as discussed in Golden-Marx et al. (2023) and Huang et al. (2022), the strongest correlation between stellar mass and halo mass has been measured when the core of the BCG is excluded. Here, we investigate the inclusion of M14 and use 5 non-overlapping radial apertures, 0-30 kpc, 30 kpc-80 kpc (representative of the ICL transition region (Zhang et al., 2024)), 80 kpc-150 kpc, 150 kpc-225 kpc, 225 kpc-300 kpc. We note that the data set used in this part of the analysis is a slightly different sample as what was used in Section 5.1 because we require that the aperture have a non-zero stellar mass in each of the 5 apertures, so there are some where we can measure a non-zero total light profile (starting from the core), but not an individual aperture mass (i.e., there are some clusters where the data may not be deep enough to allow for measurements of the large outer apertures).
@@
In Figure 3, we plot the the data used to measure the SMHM relation, the stellar mass vs. the halo mass coloured using M14. Although the stellar mass range contained within each aperture varies, we fix the y-axis, to highlight the changing value. In each subplot, we overlay the median values of stellar mass and halo mass for the three previously defined M14 bins. The stellar mass-M14 correlation persists strongly in both the 0-30 kpc and 30 kpc-80 kpc apertures. Beyond this aperture, both the overall distribution and any trend with M14 becomes dominated by the noise, as demonstrated by the solid lines which begin to overlap.
@@
The trend measured for is nearly identical to what is found in Golden-Marx et al. (2023), where is significantly lower when the light from the core of the BCG is included. Since BCG cores can be treated as a standard candle (e.g., Lauer et al., 2007), this is unsurprising. However, we remind the reader that our estimate for is an upper limit because we assume the same uncertainty on the stellar mass for each of the radial bins, calibrated for the BCG 0-30 kpc region. If we are underestimating the uncertainty in the stellar mass of the ICL, then we we would be overestimating .
Lastly, the most interesting and surprising result comes from , which is statistically more than 5 from 0 for the 0-30 kpc, 30 kpc-80 kpc, and 80 kpc-150 kpc apertures. This means that the stellar mass-M14 correlation persists out to 150 kpc, though its strength decreases drastically beyond 80 kpc. Moreover, the value of is statistically equivalent in the BCG’s core and the BCG+ICL transition region (30 kpc-80 kpc). Since M14 is independent of the light from this radial regime, this may offer further support that the light in the transition region is also growing as a result of BCG hierarchical assembly. Additionally, we note that at large radii (r > 150 kpc), we detect no correlation between stellar mass and M14, which may lead to the decrease in measured when the BCG’s core is included and shows that for that dataset , the correlation between M14 and stellar mass is coming from the inner region (between 0-80 kpc or 0-150 kpc) of the BCG+ICL system.
@@ 6.1 The SMHM Relation
Figures 1 and 3 illustrate the connection between M14 and the ICL as inferred from the SMHM relation. Table 3 and Figure 4 quantify that the ICL and M14 correlate, such that for certain radial bins, the inclusion of M14 yields a tighter SMHM relation. However, as shown in Figure 4, the strength of this correlation weakens as we extend the ICL out to larger apertures and stops being statistically significant at radii beyond 150 kpc.
@@
For this portion of the analysis, we stack the measurements of the ICL in 5 M14 bins (20% quintiles), and then measure the mean ICL fraction and M14 in each bin. For this analysis, we do not further bin the data by halo mass as Zhang et al. (2024), using the same ICL measurements, detect no correlation between ICL fraction and richness. Moreover, we note that there is no correlation between M14 and halo mass in our measurements.
@@
As shown in Figure 5, we for the first time using observations find that clusters characterized by a large M14 have significantly larger ICL fractions. Each ICL fraction increases by between 50-75% when comparing low M14 (M14 < 1.0) and high M14 (M14 > 2.0) clusters. As a result of hierarchical merging, we expect a fraction of stars will be ejected into the ICL while another fraction will increase the BCG stellar mass. Thus, the M14 and ICL fraction correlation directly results from BCG hierarchical assembly. Therefore, it follows the increase in the fraction as a function of M14 is much larger when the BCG’s light is included (comparing the middle and lower panels), as the BCG’s stellar mass increases due to hierarchical assembly. Moreover, following the interpretation of M14 from Golden-Marx & Miller (2018) and M12 from Vitorelli et al. (2018) – clusters with larger M14 values form earlier and have had more time to evolve – we posit that this may mean that clusters characterized by larger ICL fractions are dynamically older and that as clusters evolve, the BCG+ICL system (and as a result M14 and ICL fraction) grows. Additionally, though not shown, similar to Zhang et al. (2024), we find that the ICL fraction is dependent on aperture and decreases as the ICL is extended out to 300 kpc
@@
Previous work by Jiménez-Teja et al. (2018), Dupke et al. (2022), and Jiménez-Teja et al. (2024) have found that the ICL fraction has an inverse correlation with the dynamical state of the cluster, such that relaxed clusters are characterized by low ICL fractions, while merging clusters are characterized by larger ICL fractions. Moreover, studying fossil group galaxies, Zarattini et al. (2015) suggest that systems characterized by larger M12 values are dynamically relaxed. Additionally, in cosmological hydrodynamic simulations Yoo et al. (2024) find that clusters with larger M12 values are more relaxed. Furthermore, Yoo et al. (2024) also find a strong correlation between M12 and BCG+ICL fraction, similar to what is shown in the middle panel of Figure 5. Although we use M14, as shown in Golden-Marx & Miller (2018), there is no statistical difference between these parameters due to the measurement uncertainty. Thus, the results shown in Figure 5 are in excellent agreement with the simulations from Yoo et al. (2024). Lastly, recent observational work by Casas et al. (2024) also find that applying a high M14 criteria (M14 > 1.25) better identifies dynamically relaxed clusters.
@@
This literature comparison is not a perfect. Jiménez-Teja et al. (2018), Dupke et al. (2022), and Jiménez-Teja et al. (2024) use X-ray measurements, measure the ICL using CICLE (e.g., Jiménez-Teja et al., 2018), and do not estimate the ICL fraction within a fixed aperture, Zarattini et al. (2015) focus on lower halo mass fossil galaxies, Yoo et al. (2024) use simulations the majority of which are galaxy groups, and Casas et al. (2024) also analyze an X-ray sample and do not consider the ICL. However, the excellent agreement with the simulations from Yoo et al. (2024) suggests that both ICL fraction and M14 may correlate with dynamical state such that larger M14 values and higher ICL fractions are typical of dynamically relaxed systems. However, given this contradiction to results from Jiménez-Teja et al. (2018), Dupke et al. (2022), and Jiménez-Teja et al. (2024) we posit two alternatives. First, X-ray relaxation measurements are made on large cluster scale observations. However, M14 depends only on the cluster’s central region. Therefore, it is possible that while M14 traces the hierarchical assembly of the BCG, it is unrelated to the cluster’s overall dynamical state because it is characteristic of just the mergers within the core. Second, Zhang et al. (2024) find that the ICL fraction is dependent on the choice of radial aperture, and thus it is possible that the measurements from Jiménez-Teja et al. (2018), Dupke et al. (2022), and Jiménez-Teja et al. (2024) are not directly comparable to our own. However, further investigation, using simulations and consistent ICL measurement methods and definitions are required to gain greater insight into this apparent discrepancy.

### v2

@@ 4 The Statistical Modelling of the SMHM Relation
where N refers to a normal distribution, is the mathematical offset, is the slope of the SMHM relation, is the M14 parameter introduced in Golden-Marx & Miller (2018), and is the intrinsic scatter in stellar mass at fixed . We note that measures the slope between the stellar mass and M14. Physically, is related to hierarchical growth, such that a non-zero means that the associated stellar mass is growing at least partially hierarchically. Also, the significance of this detection is given by how many standard deviations the value of is from 0. This is a nested model, if =0, Equation 4 reverts to Equation 5. The only differences between the model used in this analysis and those in Golden-Marx & Miller (2019) and Golden-Marx et al. (2023) are that we do not account for redshift evolution and we include , respectively.
A Bayesian formalism works by convolving prior information associated with a selected model with the likelihood of the observations given that model, to yield the posterior distribution for the non-nuisance parameters in our model. To determine the posterior distributions for each SMHM relation parameter, our MCMC model generates values for the observed aperture stellar masses, richness-estimated halo masses, and M14s at each step in our likelihood analysis, which are directly compared to the observed measurements.
We model the log10 BCG stellar masses (), log10 halo masses (), and M14 values () as normal distributions with mean values (locations) taken from our measurements. The standard deviations associated with each value is taken from measurement uncertainties, described in Section 3 (, , ) as well as a stochastic component from a beta function (Beta(0.5,100)) (Golden-Marx & Miller, 2018), which allows for additional uncertainty on the observational errors. These errors are treated statistically in the Bayesian model as free nuisance parameters , , and . Additionally, as done in Golden-Marx & Miller (2019), Golden-Marx et al. (2022), and Golden-Marx et al. (2023), we reduce the covariance between and by subtracting off the median values of the upper and lower limits in log/ and log/, 11.3 and 14.55, respectively. For consistency we subtract off the median stellar mass measured within 0-30 kpc in each aperture bin.
@@
The parameters of the SMHM relation are given in Table 2, where each cluster is a component in the summed log likelihood and the terms marked are representative of the observed data. We express the entire posterior as:
@@ 5 The Observed SMHM Relation
As discussed in Section 1, many recent analyses have focused on the connection between the ICL and the cluster’s dark matter halo. Recent observations (e.g., Montes & Trujillo, 2019) suggest the light distribution of the ICL may trace the halo’s dark matter distribution. Moreover, prior results have identified statistically significant correlations between the stellar mass contained within the ICL and the cluster’s halo mass (e.g., Sampaio-Santos et al., 2021; Huang et al., 2022; Golden-Marx et al., 2023; Kwiecien et al., 2024). As presented in Golden-Marx et al. (2023), accounting for the ICL increases the slope, , of the SMHM relation by 0.1-0.15. Thus, as a result of BCG+ICL hierarchical growth, the larger slope is measured because ICL, due to recent growth, is more correlated with the dark matter halo mass. Here, for the first time, we analyse the correlation between M14 and the ICL and investigate both quantitatively and qualitatively the impact of incorporating the ICL and M14 into the SMHM relation.
@@ 5.1 The BCG+ICL+M14 SMHM relation
Figure 3 visualizes the stellar mass within the selected aperture plotted against the halo mass for each annulus, colour coded by M14. In each subplot, we overplot three lines, which are representative of the median of one-third of the data sorted by M14. The median M14 values are 0.90 (dark blue), 1.50 (green) and 2.10 (yellow). Based on Equation 4, not accounting for noise, at fixed halo mass, the separation of these lines is given by , where 0.6 is the difference in the average M14 values. Although not shown, we note that the 1 values for these median lines increases with aperture size from 0.1 to 0.2. In the final subplot, we overlay the fits made to Equation 4 for our observed data as described by the posterior distribution from our MCMC analysis. For example, we fix M14 to the upper and lower M14 values to illustrate our fitting.
@@
To quantify the impact of including the ICL and M14 on the parameters of the SMHM relation, we present the median and 1 values for each of the SMHM relation parameters measured from the posterior distribution (both for when M14 is and is not incorporated) in Table 4. Additionally, we present an example 2D posterior distribution in Figure 8 in the Appendix. For consistency, we see that the results in the radial range of kpc are in excellent agreement with the results presented in Golden-Marx & Miller (2019) and Golden-Marx et al. (2022). Moreover, these SMHM relation parameters are within 1 of the values of and measured in Golden-Marx et al. (2023) (due to the different median subtraction values, we can’t compare ), which shows our results are not biased by the use of richness-inferred halo masses as opposed to SZ-estimated halo masses.
@@
Across each aperture, we see clear trends such that the incorporation of M14 increases by 0.10-0.15 matching prior results (Golden-Marx & Miller, 2018, 2019; Golden-Marx et al., 2022). Moreover, reaches an asymptote at a radius of 150 kpc, in strong agreement with the trends found in Golden-Marx & Miller (2019) and Golden-Marx et al. (2023). Additionally, the inclusion of the ICL increases by 0.1-0.15 as found in Golden-Marx et al. (2023) (a similar increase to what is found by including M14). As in Golden-Marx & Miller (2019), the value asymptotes at is significantly higher when M14 is accounted for, by approximately 0.1.
@@
Additionally, although Figure 4 does not present the measurements, the median and 1 values are given in Table 4 and range from 38 to 13 from 0. Thus, the M14-stellar mass correlation persists out to 300 kpc. However, decreases as we progress to larger radii, which is shown by the slightly decreasing separation between the solid lines shown in Figure 3.
@@
The presence of a non-zero along with the reduction in and increase in allow us to conclude that M14 is a latent parameter when the SMHM relation is measured out to the ICL. For further verification, we randomly reassign the M14 values. If a non-zero value is measured for , this would suggest an intrinsic, non-physical, correlation between these parameters. However, as shown in Table 7, for each bin, the shuffled value is equivalent to 0. Moreover, the values of , , and are in excellent agreement between the versions with the randomly assigned M14 (Table 7) and when M14 is not incorporated (Table 4). Thus, any change in parameter measurements results solely from a correlation between the BCG+ICL stellar mass and M14. Going forward, we include M14 in all measurements discussed in this analysis (though we provide the posteriors when M14 is not accounted for in Tables 4, 5, and 6).
5.2 Comparison of the SMHM relation of the BCG to the SMHM relation of outskirts
@@
Although the correlation between the stellar mass and the halo mass increases as the BCG+ICL light profile is extended outward, as discussed in Golden-Marx et al. (2023), Huang et al. (2022), and Kwiecien et al. (2024), the strongest correlation between stellar mass and halo mass has been measured when the core of the BCG is excluded. Here, we investigate the inclusion of M14 and use 5 non-overlapping radial apertures, 0-30 kpc, 30 kpc-80 kpc (representative of the ICL transition region (Zhang et al., 2024)), 80 kpc-150 kpc, 150 kpc-225 kpc, 225 kpc-300 kpc. We note that the data set used in this part of the analysis is a slightly different sample as what was used in Section 5.1 because we require that the aperture have a non-zero stellar mass in each of the 5 apertures, so there are some clusters where we can measure a non-zero total light profile (starting from the core), but not an individual aperture mass (i.e., there are some clusters where the data may not be deep enough to allow for measurements of the large outer apertures).
@@
In Figure 5, we plot the data used to measure the SMHM relation, the stellar mass vs. the halo mass coloured using M14. Although the stellar mass range contained within each aperture varies, we fix the y-axis, to highlight the changing value. In each subplot, we overlay the median values of stellar mass and halo mass for the three previously defined M14 bins. Although not shown, the 1 value increases from 0.1 to 0.3 dex as the aperture moves outward. In the final subplot, we again overlay the fits made to Equation 4 for our observed data, which is given by the posterior distribution from our MCMC analysis, where we fix M14 to the upper and lower values to illustrate our fitting. The stellar mass-M14 correlation persists strongly in both the 0-30 kpc and 30 kpc-80 kpc apertures. Beyond this aperture, both the overall distribution and any trend with M14 becomes dominated by the noise, as demonstrated by the solid lines which begin to overlap.
@@
The trend measured for is nearly identical to what is found in Golden-Marx et al. (2023), where is significantly lower when the light from the core of the BCG is included. Since BCG cores can be treated as a standard candle (e.g., Lauer et al., 2007), this is unsurprising. However, we remind the reader that our estimate for is an upper limit because we assume the same uncertainty on the stellar mass for each of the radial bins, calibrated for the BCG 0-30 kpc region. If we are underestimating the uncertainty in the stellar mass of the ICL, then we would be overestimating .
Lastly, the most interesting and surprising result comes from , which is statistically more than 5 from 0 for the 0-30 kpc, 30 kpc-80 kpc, and 80 kpc-150 kpc apertures (the maximum value is 38 and then decreases to 5 as we move to farther out apertures). This means that the stellar mass-M14 correlation persists out to 150 kpc, though its strength decreases drastically beyond 80 kpc. Moreover, the value of is statistically equivalent in the BCG’s core and the BCG+ICL transition region (30 kpc-80 kpc). Since M14 is independent of the light from this radial regime, this may offer further support that the light in the transition region is also growing as a result of BCG hierarchical assembly. Additionally, we note that at large radii (r > 150 kpc), we detect no correlation between stellar mass and M14 (the significance of is less than 2), which may lead to the decrease in measured when the BCG’s core is included and shows that for that dataset, the correlation between M14 and stellar mass is coming from the inner region (between 0-80 kpc or 0-150 kpc) of the BCG+ICL system.
@@ 6.1 The SMHM Relation
Figures 3 and 5 illustrate the connection between M14 and the ICL as inferred from the SMHM relation. Table 4 and Figure 6 quantify that the ICL and M14 correlate, such that for certain radial bins, the inclusion of M14 yields a tighter SMHM relation. However, as shown in Figure 6, the strength of this correlation weakens as we extend the ICL out to larger apertures and stops being statistically significant at radii beyond 150 kpc.
@@
For this portion of the analysis, we stack the measurements of the ICL in 5 M14 bins (20% quintiles), and then measure the mean ICL fraction and M14 in each bin. For this analysis, we do not further bin the data by halo mass as Zhang et al. (2024), using the same ICL measurements, detect no correlation between ICL fraction and richness. This lack of correlation between ICL fraction and halo mass is also found in observations from Ragusa et al. (2023) and simulations presented in Contini et al. (2014), Brough et al. (2024), and Contreras-Santos et al. (2024). Moreover, we remind the reader that Figure 2 shows that there is no correlation between M14 and halo mass in our measurements since the median value is constant across the entire halo mass range. Additionally, we note that the median halo mass differs by less than 1 within each of the M14 bins.
@@
As shown in Figure 7, we for the first time using observations find that clusters characterized by a large M14 have significantly larger ICL fractions. Each ICL fraction increases by between 50-75% when comparing low M14 (M14 < 1.0) and high M14 (M14 > 2.0) clusters. As a result of hierarchical merging, we expect a fraction of stars will be ejected into the ICL while another fraction will increase the BCG stellar mass. Thus, the M14 and ICL fraction correlation directly results from BCG hierarchical assembly. Therefore, it follows that the increase in the fraction as a function of M14 is much larger when the BCG’s light is included (comparing the middle and lower panels), as the BCG’s stellar mass increases due to hierarchical assembly. Moreover, following the interpretation of M14 from Golden-Marx & Miller (2018) and M12 from Vitorelli et al. (2018) – clusters with larger M14 values form earlier and have had more time to evolve – we posit that this may mean that clusters characterized by larger ICL fractions are dynamically older and that as clusters evolve, the BCG+ICL system (and as a result M14 and ICL fraction) grows. Additionally, though not shown, similar to Zhang et al. (2024), we find that the ICL fraction is dependent on the choice of aperture and decreases as the ICL is extended out to 300 kpc
Although early works (Rudick et al., 2006; Contini et al., 2014; Cui et al., 2014) showed evidence of a correlation between the ICL fraction and the cluster’s dynamical state, this has become a topic of focus in many recent works in both simulations and small observational samples. Using semi-analytic models, Contini et al. (2023) show that the ICL fraction positively correlates with halo concentration, a tracer of cluster dynamical state, such that more relaxed clusters have larger ICL fractions. Similarly, Brough et al. (2024), using the Magneticum Pathfinder simulation, find that the ICL fraction negatively correlates with the mass fraction of the eighth subhalo, another tracer of cluster dynamical state, such that more relaxed clusters have larger BCG+ICL fractions. Furthermore, using The 300 Project, Contreras-Santos et al. (2024) find that the ICL fraction negatively correlates with both the offset of the centre of mass from the BCG and the subhalo mass fraction, both dynamical state tracers, so that more relaxed clusters form earlier and have a larger ICL fraction.
While a consensus exists in simulations and semi-analytic models, there are discrepancies in observations. Using six Hubble Frontier Field clusters Montes & Trujillo (2018) find that their most relaxed cluster has a larger ICL fraction (measured using elliptical apertures between 50 kpc and 120 kpc) than the least relaxed cluster. Similarly, Ragusa et al. (2023) find a modest positive correlation between the ICL fraction and the early-type galaxy fraction, an observational tracer of more dynamically evolved and relaxed clusters. However, in contrast, Jiménez-Teja et al. (2018), Dupke et al. (2022), and Jiménez-Teja et al. (2024) have found the ICL fraction has an inverse correlation with the dynamical state of the cluster, such that relaxed clusters are characterized by low ICL fractions, while merging clusters are characterized by larger ICL fractions.
@@
As previously noted, the magnitude gap is also viewed as a tracer for cluster dynamical state. In the context of observations of fossil group galaxies, Zarattini et al. (2015) suggest that systems characterized by larger M12 values are dynamically relaxed. Additionally, recent observational work by Casas et al. (2024) find that applying a high M14 criteria (M14 > 1.25) better identifies dynamically relaxed clusters. This trend is further supported by cosmological hydrodynamic simulations from Yoo et al. (2024), which find that clusters with larger M12 values are more relaxed. Moreover, Yoo et al. (2024) also find a strong correlation between M12 and BCG+ICL fraction, similar to what is shown in the middle panel of Figure 7. Although we use M14, as shown in Golden-Marx & Miller (2018), there is no statistical difference between these parameters due to the measurement uncertainty. Thus, the results shown in Figure 7 are in excellent agreement with the simulations from Yoo et al. (2024).
@@
This literature comparison is not perfect. For example, Brough et al. (2024) look at lower mass clusters while Zarattini et al. (2015), Ragusa et al. (2023), and Yoo et al. (2024) focus on lower halo mass groups, Montes & Trujillo (2018) have only 6 clusters, Jiménez-Teja et al. (2018), Dupke et al. (2022), and Jiménez-Teja et al. (2024) use X-ray measurements, measure the ICL using CICLE (e.g., Jiménez-Teja et al., 2018), and do not estimate the ICL fraction within a fixed aperture, and Casas et al. (2024) also analyze an X-ray sample and do not consider the ICL. However, the excellent agreement with the simulations from Yoo et al. (2024) and agreement with the previously described works strongly support that both ICL fraction and M14 correlate with dynamical state such that larger M14 values and higher ICL fractions are typical of dynamically relaxed systems. However, given this contradiction to results from Jiménez-Teja et al. (2018), Dupke et al. (2022), and Jiménez-Teja et al. (2024) we posit an explanation. Zhang et al. (2024) find that the ICL fraction is dependent on the choice of radial aperture, and thus it is possible that the measurements from Jiménez-Teja et al. (2018), Dupke et al. (2022), and Jiménez-Teja et al. (2024) are not directly comparable to our own. However, further investigation, using simulations and consistent ICL measurement methods and definitions are required to gain greater insight into this apparent discrepancy.